# Sea-Land-Cloud Segmentation in Satellite Hyperspectral Imagery by Deep Learning

### v2

@@ 2.4 Segmentation Performance Metrics and
2.4.1 Proposed On-Board Sea-Land-Cloud Ranking System
@@
Before introducing the metrics to evaluate the segmentation outcomes, we next introduce example applications to study the utility of the sea, land, and cloud segments in increasing the satellite’s autonomy, while reducing data postprocessing and decision-making latency. Several performance metrics discussed in the following section are indeed based on the proposed example applications. Namely, in Fig. 2 we suggest a novel sea-land-cloud ranking system intended for in-orbit deployment. The system prioritizes HS images for downlink based on sea, land, and cloud coverage levels, ensuring high-priority transmission. The images are arranged in ascending order of cloud coverage levels, descending order of sea coverage levels, and also decreasing order of land coverage levels, with the first positions reserved for images of higher downlink priority. Unlike existing ranking systems that target only clouds [13], our proposed approach is the first of its kind for HSI-equipped satellites, simultaneously locating and ranking sea, land, and clouds within priority queues. In Fig. 2, we also suggest other on-board use cases that further highlight the usefulness of the sea-land-cloud segments. The use cases include alternatives such as discarding or applying selective lossy compression to non-essential pixels to optimize the utilization of the Shannon channel capacity, particularly vital for satellites like HYPSO-1 that may be operating with limited line-of-sight (LOS) towards ground stations [29].
@@
Figure 2: An intelligent on-board processing pipeline integrating a ranking system based on sea-land-cloud segmentation. $T\!H_{cl}$, $T\!H_{sea}$, and $T\!H_{land}$ represent thresholds configured at the ground center to regulate data management during in-flight operations.
2.4.2 Performance Metrics
@@
Regarding performance metrics, we utilize a range of metrics to assess the quality of the segmented images on the test set (prior to generalization tests in the deployment set) in order to evaluate the models’ performance after training. First, our primary focus is on average accuracy, a widely used metric for evaluating segmentation models. Unlike overall accuracy, this average metric is particularly convenient for imbalanced test sets, providing a more realistic assessment. Furthermore, adhering to the literature of cloud detection in satellite imagery [12, 14], we also calculate the false and true alarm ratios, for each class in this work. Finally, in our supplementary material supporting further the conclusions of our work, we present additional state-of-the-art metrics that we consider secondary in importance for our purpose compared to the previous metrics. Namely, the F-1 score in addition to the overall accuracy.
Furthermore, we introduce additional performance metrics for the ranking system to assist the evaluation of the ML and DL models. To assess the predicted rankings for each coverage level, we use the Spearman’s rank correlation coefficient to measure the relationship between the predicted and ground-truth rankings, with values ranging from 1-1 to 1, representing a perfect match at the upper end. Moreover, we compute the Mean Absolute Error (MAE) for each coverage level, treating each level similar to a separate regression task. To maintain simplicity, we include the details about MAE in our supplementary materials.
@@
In the context of on-board data management proposed in Fig. 2, the false alarm ratios due to misclassifications indicate either the potential loss of data or the transmission of irrelevant data during downlink, as explained next. In cloud segmentation, the True Positive Ratio (TPR) involves correctly detecting cloudy pixels, while the True Negative Ratio (TNR) relates to recognizing non-cloudy data with potential value. The False Positive Ratio (FPR) is crucial because it indicates misclassified pixels, wrongly identified as clouds but that actually correspond to sea or land data. Following Fig. 2, these missclassifications can result in the discard or selective lossy compression with lower bitrate of valuable data. On the other hand, the False Negative Ratio (FNR) is less critical than FPR, as it does not involve data loss or data degradation, but only a less efficient usage of the downlink channel transmitting undetected pixels actually corresponding to clouds. In the context of sea segmentation, true positives signify correct identification of valuable sea data, while true negatives correctly recognize non-relevant data for potential deletion or compression. In sea-focused campaigns, FNR is critical, as it deals with misclassified sea-related pixels, potentially leading to loss of valuable data or degradation. In contrast, FPR is of less significance since its impact lies merely in downlink efficiency with the transmission of irrelevant pixels (clouds or land), rather than causing any actual data loss. The same analysis and conclusion apply when examining the context of land segmentation. In short, without taking into account the earlier suggested reasonable compromise between false positives and false negatives, for both sea and land categories, FNR has priority over FPR due to valuable data loss or degradation in the case of high FNR. However, for cloud data, FPR is critical as it leads to irreversible data loss or lossy compression, with FNR affecting merely downlink efficiency.
@@
Table II: Metrics for segmentation models trained using 112 spectral channels.
1D-ML/SGD 0.91 0.90 - 72
@@
1D-Justo-LiuNet 0.93 1.00 4,563 15,136
1D-Justo-HuNet 0.91 1.00 9,543 10,966
1D-Justo-LucasCNN 0.91 0.90 25,323 11,554
2D-NU-Net-mod 0.8 0.83 32,340 364
@@
Table III: Ranking of models’ metrics from best to worst for segmentation models trained over 112 channels.
@@
1 1D-Justo-LiuNet 1D-Justo-LiuNet 1D-Justo-LiuNet 1D-ML/SGD
2 2D-Justo-UNet-Simple 1D-Justo-HuNet 2D-Justo-UNet-Simple 1D-ML/LDA
3 1D-ML/SGD 2D-Justo-UNet-Simple 1D-Justo-HuNet 2D-CUNet Reduced
4 1D-Justo-HuNet 1D-ML/LDA 2D-CUNet++ Reduced 2D-CUNet
5 1D-Justo-LucasCNN 1D-Justo-LucasCNN 2D-CUNet Reduced 2D-CUNet++
6 2D-CUNet++ Reduced 1D-ML/SGD 2D-CUNet++ 2D-CUNet++ Reduced
7 1D-ML/LDA 2D-CUNet Reduced 1D-Justo-LucasCNN 2D-Justo-UNet-Simple
8 1D-ML/QDA 2D-UNet FAUBAI 2D-NU-Net-mod 2D-NU-Net-mod
9 2D-UNet FAUBAI 2D-CUNet 2D-NU-Net 2D-NU-Net
10 2D-UNet FAUBAI Reduced 2D-NU-Net-mod 2D-CUNet 2D-UNet FAUBAI Reduced
11 2D-CUNet++ 2D-CUNet++ Reduced 2D-UNet FAUBAI Reduced 2D-UNet FAUBAI
12 2D-NU-Net 2D-UNet FAUBAI Reduced 2D-UNet FAUBAI 1D-ML/NB
13 2D-NU-Net-mod 2D-CUNet++ - 1D-ML/QDA
14 2D-CUNet Reduced 2D-NU-Net - 1D-Justo-HuNet
15 2D-CUNet 1D-ML/QDA - 1D-Justo-LucasCNN
16 1D-ML/NB 1D-ML/NB - 1D-Justo-LiuNet
Figure 3: Comparison of false cloud detection ratios for segmentation models trained over 112 channels.
Figure 4: Comparison of false sea detection ratios for segmentation models trained over 112 channels.
Figure 5: Comparison of false land detection ratios for segmentation models trained over 112 channels.

### v3

@@ II-D Segmentation Performance Metrics and
II-D1 Proposed On-Board Sea-Land-Cloud Ranking System
@@
Before introducing the metrics to evaluate the segmentation outcomes, we next introduce example applications to study the utility of the sea, land, and cloud segments in increasing the satellite’s autonomy, while optimizing operations, and reducing decision-making latency. Several performance metrics discussed in the following section are based on the next proposed example applications. In Fig. 2, we suggest a novel sea-land-cloud ranking system for in-orbit deployment. The system prioritizes HS images for downlink based on sea, land, and cloud coverage levels, ensuring high-priority transmission. The images are arranged in ascending order of cloud coverage levels, descending order of sea coverage levels, and also decreasing order of land coverage levels, with the first positions reserved for images of higher downlink priority. Unlike existing ranking systems that target only clouds [13], our proposed approach is the first of its kind for HSI-equipped satellites, simultaneously locating and ranking sea, land, and clouds within priority queues. In Fig. 2, we also suggest other on-board use cases that further highlight the usefulness of the sea-land-cloud segments. The use cases include alternatives such as discarding or applying selective lossy compression (e.g., via Compressive Sensing [23]) to non-essential pixels to optimize the utilization of the Shannon channel capacity, particularly vital for satellites such as HYPSO-1, which have constrained downlink bandwidth and limited line-of-sight (LOS) to ground stations [29].
@@
Figure 2: An automated on-board processing pipeline integrating a ranking system based on sea-land-cloud segmentation. $T\!H_{cl}$, $T\!H_{sea}$, and $T\!H_{land}$ represent thresholds configured at the ground center to regulate data management during in-flight operations.
II-D2 Performance Metrics
@@
Regarding performance metrics, we utilize a range of metrics to assess the quality of the segmented images on the test set (prior to generalization tests in the deployment set) in order to evaluate the models after training. The employed metrics, including accuracy [13], alarm ratios [11, 14], precision [13, 15], recall and F-1 score [13, 15], are frequent in the field of on-board segmentation for satellite spectral missions, enabling direct comparisons. However, our primary focus is on average pixel-level accuracy, a widely used metric for evaluating segmentation models. Unlike overall accuracy, this average metric is particularly convenient for imbalanced test sets, providing a more realistic assessment. Furthermore, we concentrate on the false alarm ratios for each individual class. Moreover, to evaluate the predicted rankings for each coverage level, we use the Spearman’s rank correlation coefficient, following the literature [13]. This coefficient measures the relationship between the predicted and ground-truth rankings, with values ranging from -1 to 1, representing a perfect match at the upper end. To maintain a focused analysis, we employ these concrete metrics to compare model performance. However, we note that in our supplementary material, we also provide additional metrics such as the precision, recall, F1 score, overall accuracy, and the Mean Absolute Error (MAE) for each coverage level, treating each level similar to a regression task.
@@
Regarding alarm ratios in the context of the on-board data management proposed in Fig. 2, the false alarm ratios due to misclassifications indicate either the potential loss of data or the transmission of irrelevant data during downlink, as explained next. In cloud segmentation, the True Positive Ratio (TPR) involves correctly detecting cloudy pixels, while the True Negative Ratio (TNR) relates to recognizing non-cloudy data with potential value. On the one hand, the False Positive Ratio (FPR) is crucial because it indicates misclassified pixels, wrongly identified as clouds but that actually correspond to sea or land data. Following Fig. 2, these missclassifications can result in the discard or selective lossy compression with lower bitrate of valuable data. On the other hand, the False Negative Ratio (FNR) is less critical than FPR, as it does not involve data loss or data degradation, but only a less efficient usage of the downlink channel transmitting undetected cloud pixels. In the context of sea segmentation, true positives signify correct identification of valuable sea data, while true negatives correctly recognize non-relevant data for potential deletion or compression. In sea-focused campaigns, FNR is critical, as it deals with misclassified sea-related pixels, potentially leading to loss of valuable sea data or its degradation. In contrast, FPR is of less significance since its impact lies merely in downlink efficiency with the transmission of irrelevant pixels (clouds or land), rather than causing any data loss. The same analysis and conclusion apply when examining the context of land segmentation.
In summary, without taking into account the earlier suggested reasonable compromise between false positives and false negatives, for both sea and land categories, FNR has priority over FPR due to valuable data loss or degradation in the case of high FNR. However, for cloud data, FPR is critical as it leads to irreversible data loss or lossy compression, with FNR affecting merely downlink efficiency. It is noted that data loss in any communication channel, including a satellite radio channel, is undoubtedly more critical than utilizing the channel less efficiently. This is because lost data leads to permanent loss of information, or irreversible degradation below Shannon’s compression entropy, while the inefficient use of the downlink impacts only other channel parameters such as transmission speed, transmission power and radio spectra utilization. While these downlink parameters can certainly be bottlenecks for (near) real-time applications and impact the satellite’s power budget, they do not lead to a critical scenario of absolute irrevocable data loss. Therefore, in our work, parameters affecting the downlink efficiency are given less weight compared to the irreversible loss or degradation of valuable data.
III Results
III-A Main Results
We conduct model training using 112 channels of both raw sensor data and L1b calibrated radiance, and we discuss the segmentation of raw and L1b radiance in Section IV. For L1b radiance, we present the outcomes of the segmentation models in Table II, where we utilize bold font to mark our models. The table displays the performance of 1D ML, 1D-CNNs, 2D-CNNs, and FastViT for segmentation, with the primary metric being the average accuracy across the sea, land, and cloud classes. Additionally, to assess the sea-land-cloud segmentation performance for the proposed on-board ranking system, we also include the average Spearman’s rank coefficient, in Table II, calculated across the respective rankings based on sea, land, and cloud coverage. Moreover, we also include information about the network’s parameter count to offer insights of the potential memory demands for in-orbit deployment. In addition, we provide the average inference time per image measured in a testing setup on ground, on a NVIDIA RTX A4000 GPU (for DL), and on the AMD Ryzen 7 5800X CPU (for ML). From the numbers presented in Table II, we sort the top 10 models in Table III based on their respective metrics. In Table III, we first sort the models solely based on their average accuracy, from highest to lowest, and subsequently, we sort them in the following columns, first by the Spearman’s coefficient (highest to lowest), then by the number of parameters (lowest to highest), and finally by inference time (shorter to longer). Therefore, the first rows in the table represent the top models in the test set for each respective metric.
@@ III-A Main Results
TABLE II: Metrics for segmentation models trained over 112 spectral channels.
Type Model Accuracy Spearman’s Coefficient Parameter Count Inference Time [ms]
1D 1D-ML/SGD 0.91 0.90 - 72
@@
1D-Justo-LiuNet 0.93 1.00 4,563 166
1D-Justo-HuNet 0.91 1.00 9,543 69
1D-Justo-LucasCNN 0.91 0.90 25,323 142
2D 2D-NU-Net-mod 0.80 0.83 32,340 364
@@
FastViT-T8 0.82 - 5,033,795 132
FastViT-S12 0.90 - 10,242,819 156
FastViT-MA36 0.88 - 44,647,755 288
TABLE III: Ranking of top 10 segmentation models trained over 112 channels by various metrics, from best to worst.
@@
1 1D-Justo-LiuNet 1D-Justo-LiuNet 1D-Justo-LiuNet 1D-Justo-HuNet
2 2D-Justo-UNet-Simple 1D-Justo-HuNet 2D-Justo-UNet-Simple 1D-ML/SGD
3 1D-ML/SGD 2D-Justo-UNet-Simple 1D-Justo-HuNet FastViT-T8
4 1D-Justo-HuNet 1D-ML/LDA 2D-CUNet++ Reduced 1D-Justo-LucasCNN
5 1D-Justo-LucasCNN 1D-Justo-LucasCNN 2D-CUNet Reduced FastViT-S12
6 2D-CUNet++ Reduced 1D-ML/SGD 2D-CUNet++ 1D-Justo-LiuNet
7 FastViT-S12 2D-CUNet Reduced 1D-Justo-LucasCNN 1D-ML/LDA
8 1D-ML/LDA 2D-UNet FAUBAI 2D-NU-Net-mod 2D-CUNet Reduced
9 1D-ML/QDA 2D-CUNet 2D-NU-Net 2D-CUNet
10 2D-UNet FAUBAI 2D-NU-Net-mod 2D-CUNet FastViT-MA36
While average accuracy offers a general overview on segmentation performance across the sea, land, and cloud classes, it may obscure poor model performance in specific classes. To address this, we evaluate individual class performance by presenting their respective binary false detection ratios. Namely, in Figs. 3-5, we provide a comparative visual representation of false alarm ratios for the binary segmentation of each distinct class. We note that the top-performing models get positioned in the lower-left corners, indicating the optimal trade-off between low false positives and low false negatives. Assuming the case example of the on-board ranking system, the graphs illustrate on the Y-axis the critical false alarm ratio corresponding to the loss (or degradation) of valuable data, while the X-axis shows the false alarm ratio that reflects the inefficient utilization of the downlink channel. Both target variables should be minimized. As previously described, data loss is measured as FNR for sea and land categories and as FPR for clouds. Additionally, the inefficient utilization of the channel is measured as FPR for sea and land categories and as FNR for clouds. This clarifies that in Fig. 3 for clouds, the Y axis represents FPR, and the X axis represents FNR. In Figs. 4-5 for sea and land, the Y axis represents FNR, and the X axis represents FPR. In relation to this trade-off, we stress that while minimizing the loss of valuable data remains crucial, we also highlight the importance of maintaining an acceptable downlink efficiency.
Figure 3: Comparison of false detection ratios for the cloud class by segmentation models trained over 112 channels.
Figure 4: Comparison of false detection ratios for the sea class by segmentation models trained over 112 channels.
Figure 5: Comparison of false detection ratios for the land class by segmentation models trained over 112 channels.
Moreover, Table IV shows the results after retraining various models using the 3-channel HYPSO-1 dataset. We employ this dataset to assess the impact of channel reduction on performance, parameter count, and inference time. In addition, the results in Table IV are further confirmed by training once again all models with a distinct set of three channels different from the bands used in Section II-B. The new results align with the previous findings and for simplicity, we include them in the supplementary material.